y is directly proportional to x, and y = 20 when x = 6. What is y when x = 14?
Answer by Alan3354(69443)

y is directly proportional to x, and y = 20 when x = 6. What is y when x = 14?
----------
y = k*x
20 = k*6
k = 10/3
==================
What is y when x = 14?
y = 14*10/3 = 140/3
--------------------
You don't have to solve for k:
20/6 = y/14
Cross multiply
6y = 280
y = 140/3